OpenStudy (anonymous):

I'm stuck on scientific notation.. (6.4 x 10^6) - (5.3 x 10^4)

OpenStudy (anonymous):

that is the same as (640 x 10^4) - (5.3 x 10^4)

OpenStudy (anonymous):

(640 - 5.3) x 10^4 = 634.7 x 10^4 = 6.347 x 10^2 x 10^4 = 6.347 x 10^6

OpenStudy (anonymous):

Okay, so would that mean 5.3 is the same as 530?

OpenStudy (anonymous):

Thanks!